OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (3, -3) with 0° ≤ θ < 360°.

OpenStudy (jdoe0001):

$$ \text{from (3,-3) we can see that x=3, y=-3; THEN}\\ \text{you want to get the polar coords of (radius, angle) pair}\\ \text{using pythagorean identity } r^2=x^2+y^2\\ r=\sqrt{(3)^2+(-3)^2} \implies r=\sqrt{18} \implies \boldsymbol{r=2\sqrt{3}}\\ \text{now to find the angle } tan(\theta)=\cfrac{y}{x} \\ tan(\theta)=\cfrac{-3}{3}=-1 \implies tan(tan\theta)^{-1}=tan(-1)^{-1} $$

OpenStudy (jdoe0001):

can you get that angle WHOSE tangent is -1?